# Continuous-Time Value Iteration for Multi-Agent Reinforcement Learning

### v1

@@ 5 Conclusion and Limitations
We propose a novel approach that integrates PINNs into the actor-critic framework to solve CT-MARL problems. Specifically, we approximate value functions using HJB-based PINNs and introduce VGI to improve value approximations, thus mitigating the adverse impact of inaccurate value approximations on policy learning. We validate our VIP across continuous-time variants of MPE and MuJoCo environments and empirically demonstrate that VIP converges faster and achieves higher accumulated reward compared to baselines of SOTA. Furthermore, we investigate the importance of activation function choice and loss term weighting for VIP performance. In summary, our proposed VIP offers a promising approach to solve CT-MARL problems. However, we notice the limitation of VIP: the current work only focuses on cooperative CT-MARL tasks. Many real-world multi-agent scenarios involve competitive interactions, which have not yet been explored within CT-MARL frameworks. In competitive settings, agents no longer share a common value function, and such values often become discontinuous due to collision penalties in the reward function, which poses convergence challenges for PINN Zhang et al. (2024b). As part of future work, we aim to extend VIP to explore competitive CT-MARL problems, particularly those involving state constraints.
@@ B.1 Coupled Oscillator
We evaluate on a two-agent coupled spring–damper system. Each agent $i\in\{1,2\}$ controls one mass in a pair of identical oscillators with linear coupling. The continuous-time dynamics are
@@
we can compute the exact infinite-horizon LQR solution:
@@
    Form the optimal state-feedback gain
@@ B.2 Continuous-Time MPE
We build on the standard MPE of Lowe et al. [2017b], which simulates $N$ holonomic agents in a 2D world with simple pairwise interactions. In the original MPE each control step advances the physics by a fixed time-step $\Delta t_{\rm fixed}=0.1\,$s. To evaluate our continuous-time framework under irregular sampling, we modify the simulator so that at each step the integration interval is drawn randomly,
@@ B.3 Continuous-Time Multi-Agent MuJoCo
For high-dimensional control we adapt the discrete-time Multi-Agent MuJoCo suite (e.g. cooperative locomotion, quadruped rendezvous). By default MuJoCo uses an internal physics integrator with a base time-step of 0.010.01\,s and repeats each action for $K_{\rm fixed}=5$ frames, yielding an effective control interval $\Delta t_{\rm fixed}=0.05\,$s. We instead sample the number of frames per control step,
@@ C.1 Value Gradient Comparison
Fig. 10 compares VIP’s learning curves when the policy network uses ReLU or Tanh activations on Ant $2\times 4d$ and Ant $4\times 2$. Across both tasks the Tanh implementation converges faster and attains a higher plateau reward, whereas the ReLU version peaks earlier and then undergoes a mild performance decay. The observation aligns with the earlier Tanh-versus-ReLU analysis reported in the main paper Fig. 5: smoother activation functions mitigate gradient saturation and promote more stable policy updates. The additional evidence from the two Ant variants therefore reinforces our previous claim that Tanh is better suited for value–gradient propagation in VIP.

### v2

@@ 5 Conclusion and Limitations
We propose a novel approach that integrates PINNs into the actor-critic framework to solve CT-MARL problems. Specifically, we approximate value functions using HJB-based PINNs and introduce VGI to improve value approximations, thus mitigating the adverse impact of inaccurate value approximations on policy learning. We validate our VIP across continuous‑time variants of MPE and MuJoCo environments and empirically demonstrate that VIP converges faster and achieves higher accumulated reward compared to baselines of SOTA. Furthermore, we investigate the importance of activation function choice and loss term weighting for VIP performance. In summary, our proposed VIP offers a promising approach to solve CT-MARL problems. However, we notice the limitation of VIP: the current work only focuses on cooperative CT-MARL tasks. Many real-world multi-agent scenarios involve competitive interactions, which have not yet been explored within CT-MARL frameworks. In competitive settings, agents no longer share a common value function, and such values often become discontinuous due to collision penalties in the reward function, which poses convergence challenges for PINN [51]. As part of future work, we aim to extend VIP to explore competitive CT-MARL problems, particularly those involving state constraints.
@@ B.1 Coupled Oscillator
We evaluate on a two‐agent coupled spring–damper system. Each agent $i\in\{1,2\}$ controls one mass in a pair of identical oscillators with linear coupling. The continuous‐time dynamics are
@@
we can compute the exact infinite‑horizon LQR solution:
@@
    Form the optimal state‑feedback gain
@@ B.2 Continuous-Time MPE
We build on the standard MPE of [25], which simulates $N$ holonomic agents in a 2D world with simple pairwise interactions. In the original MPE each control step advances the physics by a fixed time‐step $\Delta t_{\rm fixed}=0.1\,$s. To evaluate our continuous‐time framework under irregular sampling, we modify the simulator so that at each step the integration interval is drawn randomly,
@@ B.3 Continuous-Time Multi-Agent MuJoCo
For high‑dimensional control we adapt the discrete‐time Multi‑Agent MuJoCo suite (e.g. cooperative locomotion, quadruped rendezvous). By default MuJoCo uses an internal physics integrator with a base time‐step of 0.010.01\,s and repeats each action for $K_{\rm fixed}=5$ frames, yielding an effective control interval $\Delta t_{\rm fixed}=0.05\,$s. We instead sample the number of frames per control step,
@@ C.1 Value Gradient Comparison
Fig. 10 compares VIP’s learning curves when the policy network uses ReLU or Tanh activations on Ant $2\times 4d$ and Ant $4\times 2$. Across both tasks the Tanh implementation converges faster and attains a higher plateau reward, whereas the ReLU version peaks earlier and then undergoes a mild performance decay. The observation aligns with the earlier Tanh‐versus‐ReLU analysis reported in the main paper Fig. 5: smoother activation functions mitigate gradient saturation and promote more stable policy updates. The additional evidence from the two Ant variants therefore reinforces our previous claim that Tanh is better suited for value–gradient propagation in VIP.